OpenStudy (anonymous):

find the constants m and b in the linear function f(x)=mx+b such that f(0)=2 and f(3)=-1

OpenStudy (smored):

f(0)=0m+b=2 b=2 f(3)=m(3)+2=-1 3m=-3 m=-1

OpenStudy (anonymous):

Thank you! Been years since I have taken a math class. Wasn't sure how to set it up. I appreciate it.

OpenStudy (mathmale):

Write out the unknown function: f(x)=mx+b. Now substitute x=0 (as smored has already done). Result: f(0)=0+b. We are told that f(0)=2. Therefore, as smored has demonstrated, f(0)=2=0+b, so b=2.

OpenStudy (smored):

no prob!